OpenStudy (anonymous):

Determine algebraically whether the function is even, odd, or neither even nor odd. f(x) = -9 x3 + 8x

OpenStudy (anonymous):

both exponents are ...

OpenStudy (anonymous):

odd

OpenStudy (anonymous):

bingo

OpenStudy (tkhunny):

That is not an "algebraic" solution.

OpenStudy (anonymous):

so wait you dont't have to do that long algebraic stuff?

OpenStudy (anonymous):

on the other hand, that is not what your math teacher wants you to say

OpenStudy (anonymous):

haha true thats what I thought but she doesnt need to see my work lol

OpenStudy (anonymous):

yeah your math teacher wants you to compute \[f(-x)\] and show that it is \(-f(x)\)

OpenStudy (anonymous):

\[f(-x)=-9(-x)^3+8(-x)=\text{ some algebra}=-(-9x^3+8x)=-f(x)\]

OpenStudy (anonymous):

yeah haha I like the easy way but thanks!

OpenStudy (anonymous):

yw